A company currently pays a dividend of $4 per share (D0= $4). It is estimated that the company’s dividend will grow at a rate of 10% per year for the next 2 years, and then at a constant rate of 5% thereafter. The company’s stock has a beta of 6, the risk-free rate is 4% and the market risk premium is 2%. What is your estimate of the stock’s current price?
Answer #1

Required return=Risk free rate+Beta*Market risk premium

=4+(6*2)=16%

D1=(4*1.1)=$4.4

D2=(4.4*1.1)=$4.84

Value after year 2=(D2*Growth Rate)/(Required Return-Growth Rate)

=(4.84*1.05)/(0.16-0.05)

=$46.2

Hence current price=Future dividends*Present value of discounting factor(rate%,time period)

=4.4/1.16+4.84/1.16^2+46.2/1.16^2

=$41.72(Approx).